### v1

# Towards a Method for LLM-enabled Indoor Navigation thanks: This work was founded by European Union - Next Generation EU, in the context of The National Recovery and Resilience Plan, Investment Partenariato Esteso PE8 ”Conseguenze e sfide dell’invecchiamento”, Project Age-IT, CUP: B83C22004880006

Abstract
Indoor navigation presents unique challenges due to complex layouts, lack of GPS signals, and accessibility concerns. Existing solutions often struggle with real-time adaptability and user-specific needs. In this work, we explore the potential of a Large Language Model (LLM), i.e., ChatGPT, to generate natural, context-aware navigation instructions from indoor map images. We design and evaluate test cases across different real-world environments, analyzing the effectiveness of LLMs in interpreting spatial layouts, handling user constraints, and planning efficient routes. Our findings demonstrate the potential of LLMs for supporting personalized indoor navigation, with an average of 52% correct indications and a maximum of 62%. The results do not appear to depend on the complexity of the layout or the complexity of the expected path, but rather on the number of points of interest and the abundance of visual information, which negatively affect the performance.
Index Terms:
Inclusivity, Indoor localization, Large Language Model, AI-driven path planning, Accessibility, spatial reasoning, indoor maps.
I Introduction
Efficient indoor navigation and scene representation are essential for various applications, including robotics, smart environments, and assistive technologies. Unlike outdoor navigation, which benefits from well-established global positioning systems (GPS) and extensive mapping databases, indoor navigation presents unique challenges [1]. The complexity and variability of enclosed spaces, the absence of GPS signals, and the need for detailed real-time scene understanding make indoor navigation a demanding task [2]. Current solutions rely on advanced computational models to interpret the environment and adapt to dynamic conditions [3]. However, they often lack flexibility, accessibility, and adaptability for diverse user needs.
There is a growing need for a system capable of providing personalized, natural-language navigation assistance to address these limitations[4]. Such a system should interpret indoor maps and generate human-like, context-aware navigation instructions tailored to individual users. A system like this would also be highly beneficial for visually impaired individuals, as a detailed, natural-language description of the environment can enhance spatial awareness, helping them build a mental representation of their surroundings and navigate with greater confidence and independence. Integrating multimodal inputs, such as textual queries, voice commands, or sensor data, could provide more adaptive and accessible navigation experiences, ultimately fostering greater independence for individuals facing navigational barriers[5]. Beyond accessibility challenges, navigating unfamiliar indoor environments is another scenario where such a system would be beneficial. People visiting a new location, such as airports, shopping malls, hospitals, or office buildings, often struggle with orientation, requiring clear and accessible guidance to reach their destinations efficiently. A common approach to overcoming these challenges is to ask other people for directions, who typically provide landmark-based instructions (e.g., you will see a shoe shop, turn right, then you will encounter a phone charging area…) Given these challenges, leveraging Large Language Models (LLMs) for indoor navigation presents a promising alternative. LLMs can interpret complex multimodal inputs, reason over spatial relationships, and generate human-like natural language indications tailored to user-specific queries. This adaptability makes them well-suited for generating dynamic, human-like navigation instructions based on indoor maps[6]. In fact, recent advances in LLMs have opened new avenues for user-friendly navigation solutions for indoor and outdoor contexts, and preliminary studies have been performed leveraging LLMs to address this goal. For example, in [7], authors introduce TraveLLM, a system forpublic transit route planning. Their study compares multiple LLMs—including GPT-4, Claude 3, and Gemini—demonstrating that these models can integrate real-time data and user-specific constraints to form contextually aware navigation recommendations than conventional platforms (e.g., Google Maps). However, their study focuses on outdoor scenarios, leaving indoor navigation unexplored. Meanwhile, Zhang et al. [8] focus on improving navigation for people with visual impairments (PVI). NaviGPT integrates LiDAR-based obstacle detection, vibration feedback, and LLM-generated instructions to produce a real-time user experience. Unlike existing tools, e.g., Be My AI [^1] https://www.seeingai.com/ and Seeing AI  [^2] https://www.bemyeyes.com/blog/introducing-be-my-ai, which require multiple apps for obstacle detection, image recognition, or contextual directions, NaviGPT consolidates these capabilities into one cohesive system. This seamless integration offers continuous feedback without user distraction from frequent application switching. However, no evaluation of the system has been provided, and at the current stage, the proposal appears to mainly target outdoor environments. For indoor context, in [9], authors examine the development of geo-descriptions—textual narratives that describe the internal layout of a building. These authors combine classical GIS spatial analysis and LLM-based text generation to automate the creation of human-readable indoor maps, which is traditionally labour-intensive. However, the authors do not consider querying the system for navigation paths in the reconstructed textual representation of the building. Furthermore, the evaluation is performed on simplified cases.
Despite these preliminary studies, the use of visual maps as input for LLMs to provide indications based on user requests has not been explored. This could highly simplify the architecture of indoor navigation systems, thus making them easily adaptable to different indoor environments. Once a visual map is made available—as common for airports, shopping malls, and large environments with several points of interest—a system based on these principles is expected to be ready to use in any environment without complex, ad hoc, configurations. Indeed, this approach avoids the task of providing structured representations of the maps for each environment, which can be time-consuming.
To address the research gap, we evaluate the potential of an LLM, namely GPT-4, to assist in indoor navigation using three real-world interior 2D maps of buildings, including two airport terminal maps and one shopping mall map. We investigate the LLM ability to generate step-by-step guidance based on diverse user preferences and spatial constraints. A usage scenario is depicted in Fig. 1, where a user asks the LLM for directions to a specific gate and receives human-like indications to navigate the environment. Using different real scenarios, we want to evaluate different test cases across three real-world environments, analyzing the effectiveness of LLMs in interpreting spatial layouts, handling user constraints, and planning efficient routes.
The main contributions of this paper are as follows:
  1. 1.
    We propose an LLM-based approach for generating personalized navigation instructions based on indoor maps.
  2. 2.
    We design test cases to assess LLM performance under different conditions, such as long distances, extensive textual annotations, and complex map layouts.
  3. 3.
    We evaluate the effectiveness of GPT-4 in generating accurate, human-like navigation guidance.
By exploring the role of LLMs in indoor navigation, this work aims to bridge the gap between traditional map-based systems and intelligent, real-time, and user-adaptive guidance solutions.
Figure 1: Example of an LLM assisting with indoor navigation tasks.
All supplementary material, including prompts, input maps, and results, are publicly available  [^3] https://doi.org/10.5281/zenodo.14993277.
II Understanding user needs: Requirements for an inclusive experience
Today, many public venues have navigational aids like digital signage, online platforms, or smartphone apps. However, these aids frequently fail to accommodate individuals with different requirements, and the general public often encounters difficulties due to unclear or incorrect directions.
For instance, mobile apps that use digital maps often do not include features to assist people with visual impairments [10] or mobility disabilities. Nonetheless, navigation apps designed for accessibility have shown to be beneficial for such users [11]. A detailed and accurate description of the surrounding environment can enhance spatial awareness and improve the overall navigation experience. This is especially useful in complex indoor spaces, where clear, context-aware information can help users confidently reach their destinations. Designing navigation systems with adaptability in mind ensures they can satisfy a diverse range of needs, from visitors unfamiliar with a building’s layout to individuals who benefit from alternative navigation methods.
Indoor spaces vary widely in size, layout, and functionality, encompassing different types of destinations such as stores, restrooms, and emergency exits. Users often ask questions such as ”What can I find on this floor?” or ”Where is the nearest restroom?” highlighting the need for adaptable navigation solutions that can dynamically respond to different environments and user preferences. Among the many scenarios where such systems can be particularly beneficial is assisting users who may face additional navigation challenges, such as visual or hearing impairments, mobility constraints, or cognitive differences. While many existing systems are not designed to accommodate a broad spectrum of interaction needs, a more advanced approach could integrate personalized, multimodal assistance, ensuring that users receive navigation guidance in a format that best suits their preferences.
Figure 2: Flowgraph of the experimental setting.
Particular attention must be paid to the user experience. Navigating a complex indoor environment can be challenging for anyone, whether due to unfamiliarity with the building’s layout, unclear signage, or language barriers. These difficulties often lead users to seek assistance, emphasizing the need for a system that provides clear, accessible, and efficient guidance. To be truly effective, such a system must offer a seamless and intuitive experience.In our vision of a functional and efficient solution, users should be able to input personal preferences or specific needs and receive tailored navigation assistance in real-time. The most effective way to achieve this is through natural language interaction.
Natural language is not only a valid and immediate input method but also makes the system accessible to many users. The system can accommodate diverse needs by allowing users to interact through written or spoken text, ensuring a more inclusive and user-friendly navigation experience.
III Experimental Set-up
This section illustrates the approach we followed to evaluate the capability of the LLM to support indoor navigation. Given a user query, we expect the LLM to deliver a detailed output, including step-by-step directions, points of interest, and potential obstacles, all aimed at enhancing the user’s navigation experience.
Figure 2 describes the pipeline of our evaluation. Its four main components are:
  • Image Preprocessing Initially, we manually removed the legend from the image, to ensure that the input map is free of noise that can affect the LLM performance. This step was introduced after preliminary tests, in which we identified several errors likely caused by this information. It should be noted that this step is performed once for each used map, and not at runtime.
  • System prompt This prompt is hard-coded for the LLM and will be used for any user query. It includes all the necessary instructions, details, and potential errors to avoid. This prompt is the result of several iterations through which it has been improved to prevent critical and systematic errors.
  • User prompt The user makes a request, i.e., an LLM user prompt, aiming to navigate from point A to point B on the map.
  • User Output The system produces an output with human-like indications. This output can be delivered in written or spoken form, depending on the user preferences and possible accessibility requirements.
The model receives comprehensive contextual information by integrating the user prompt, system prompt, and sanitized map image, enabling it to generate navigational instructions. This approach ensures that the system is not only responsive to user inputs but also adaptive to various indoor environments, making it versatile across different settings.
Certain potential errors were identified and avoided during the prompt engineering phase. These errors included the model misinterpreting points that are close together on the map legend as being geographically proximate and points that appear numerically adjacent. Moreover, an overabundance of information was initially provided as output, which could overwhelm the user. A uniform approach was adopted to address these issues where a single user prompt structure—with different content—was used across all tests. This format does not vary between maps or include detailed descriptions specific to any map. Instead, references to locations in the user prompts are made either by the numerical ID of a point of interest or by the point’s name, depending on the map’s representation of the points. It should be highlighted that, in a practical scenario with a real user querying the LLM, prompt engineering is not feasible. In future work, we will study how to translate natural user queries into engineered prompts to better exploit the LLM’s potential.
The tests were conducted using three different maps: the ground floor plans of Bergamo Airport, Bologna Airport, and Orio Center—a shopping mall. These maps were sourced from the facilities’ official websites. The images differ from each other and aim to encompass various structural scenarios. The airports are characterized by a main corridor where various points of interest for guests are located on the sides, featuring many services and different types of access. Bologna Airport has a less dense set of points of interest than Bergamo Airport. In contrast, the Orio Center is characterized by a denser network of corridors, with various shops and points of interest.
For each map, five requests are made, each with a different category of difficulty. These were defined ad hoc based on a visual analysis of each map. The first three queries differ in terms of distance between starting and destination points, i.e., short (S), medium (M), and long (L). The other two queries, referred to as “specific requests” (Spec), explore other elements of complexity, which depend on each map. Specifically, we challenged the LLM with routes that have hidden or poorly marked IDs of points of interest, or with prompts in which the specific destination was not explicitly stated, e.g., ”Show me how to walk to the nearest dining area from gate number 26”. In this case, the query requires a more detailed search near gate 26 to locate the closest dining area. Table I reports examples of user prompts for the Bologna Airport.
The evaluation of the LLM’s responses is carried out by defining a ground truth of correct paths. Once the start and destination points are established for each query, we manually identify the optimal route by imagining a user standing at the starting point and then considering the shortest, least convoluted path through the structure. If there are multiple routes of similar length, they are all considered correct.
Each instruction provided by the model—i.e., the description of a path segment —is assigned a label reflecting its accuracy, which the authors decided by comparing the ground truth and the LLM output. There are two labels:
  • Correct – the instruction reflects a segment of the correct route(s) of the ground truth.
  • Incorrect – the instruction points to a different direction with respect to the destination, or asks to follow routes that are not physically possible, e.g., across walls;
Figure 3: Correct path considered for ORIO-2 query (Table  I).
Let us consider one example to exemplify the correct and incorrect cases and the rationale for our tagging. Taking the query “How to go from Piazza Italia to Milani Cafè”—cf. Figure 3 for the area of interest and the correct path—this yields the following output:
“To get to Milani Cafè from Piazza Italia, follow these steps:
  • From Piazza Italia head south towards Snipes.
  • Continue straight past Snipes, and you will see Clayton on your left.
  • Keep walking straight and pass Dainese, which will also be on your left.
  • Milani Cafè is just ahead on the right side, next to Cotton & Silk.
The green (correct) and red (incorrect) highlights are provided according to the following rationale:
  • Green directions like “continue straight past Snipes” are considered totally correct as they align with the correct path. The same is true for “keep walking straight and pass Dainese”, and the other instructions.
  • Instead, instructions such as “and you will see Clayton on your left” and “which will also be on your left” are incorrect because Clayton is actually located on the right and after Dainese. Errors related to order and left/right orientation are particularly misleading, as they may cause the user to question whether the correct direction is reversed.
It should be noted that, in practice, one incorrect indication may jeopardize the entire path. For example, users may reverse their direction based on a single incorrect instruction and subsequently get lost, even if the following directions are correct. However, in this preliminary evaluation phase, we decided to focus our evaluation on individual instructions rather than the entire path to assess the accuracy of each step independently and better isolate specific types of errors.
IV Results and Discussion
In the following, we comment on the results for each map, as each case has specific characteristics and challenges, which allow us to point out the strengths and weaknesses of the tested LLM.
IV-A Bergamo
The image of the Bergamo airport is very well-colored and clearly represented. It includes various IDs placed at specific locations on the map and additional icons that highlight points of interest, such as staircases, access points, phone charging stations, or information desks. The layout is characterized by a single corridor with a few distracting elements in the middle, and all various shops or restaurants appear along the sides. However, there are many services and points of interest to recognize, and, often, the IDs are placed very close to each other or are very small. This complicates the correct parsing of image information by the LLM.
The detailed results for Bergamo Airport are summarized in Table II. Over half of the indications are incorrect (60.78%), while 39.22%
IV-B Orio Center
The map of the Orio Center shopping mall was selected for its more sophisticated and challenging corridor layout. It features more changes in direction, and more corridors that are parallel and perpendicular to each other. Additional complexity arises from smaller retail outlets in the middle of the corridor, thus posing a challenge for the LLM to determine whether these should be considered to the right or left of the user’s path. Unlike the previous tests, all tests were conducted not with IDs to indicate the starting or ending points, but rather with the actual names of the establishments within the map, as this was the information displayed in the image.
TABLE I: User prompt examples for the Bologna Airport.
Scenario ID Query Category User Prompt Query Rationale
BOL-1 Medium Give me directions on how to walk to the Carrefour (15 on the map) from the Mondadori (7 on the map). Testing on a medium path
BOL-2 Short Give me directions on how to walk to the Vecchia Bologna Osteria (4 on the map) from the Carisbo (11 on the map). Testing on a short path.
BOL-3 Long Give me directions on how to walk to the point 2 on the map from the point 3 on the map. Testing on a long path.
BOL-4 Spec Give me directions on how to walk to the closest toilet on the map from point 3. Testing the ability to find the closest toilet and recognize the related icon.
BOL-5 Spec Give me directions on how to walk to the closest taxi station from the First Aid on the map Testing the ability to recognize difficult icons on the map and find the closest path.
Table III presents the detailed results for Orio Center. Of 58 indications, 25 were incorrect (43.10%), and 33 were correct (56.90%). This indicates an improvement in terms of performance compared to the Bergamo case. The increased intricacy of Orio Center’s layout does not negatively impact the performance of the LLM. Again, the path length does not clearly correlate with performance, as the query of category L (ORIO-3) has the highest percentage of correct indications.
TABLE II: Results for Bergamo
Query Incorrect Correct
BER-1 (S) 7 (87.50%) 1 (12.50%)
BER-2 (L) 12 (54.55%) 10 (45.45%)
BER-3 (M) 5 (62.50%) 3 (37.50%)
BER-4 (Spec) 3 (60.00%) 2 (40.00%)
BER-5 (Spec) 4 (50.00%) 4 (50.00%)
Total 31 (60.78%) 20 (39.22%)
TABLE III: Results for Orio Center
Query Incorrect Correct
ORIO-1 (M) 3 (37.50%) 5 (62.50%)
ORIO-2 (M) 9 (69.23%) 4 (30.77%)
ORIO-3 (L) 3 (18.75%) 13 (81.25%)
ORIO-4 (S) 5 (71.43%) 2 (28.57%)
ORIO-5 (L) 5 (35.71%) 9 (64.29%)
Total 25 (43.10%) 33 (56.90%)
IV-C Bologna
Similar to Bergamo airport, this map features a structure defined by a large, single corridor with shops on the sides. It differs from the first one in its use of colors, the overall representation of the floor, and a less linear configuration, with a much more pronounced central curve.
Table IV presents the detailed results for Bologna. Out of 66 indications, 27 were incorrect (40.91%), and 39 were correct (59.09%). Here, it appears that the simpler layout, with fewer points of interest, leads to better accuracy. Again, no evident relation between the query category and the percentage of incorrect instructions can be observed.
TABLE IV: Results for Bologna
Query Incorrect Correct
BOL-1 (M) 4 (28.57%) 10 (71.43%)
BOL-2 (S) 8 (47.06%) 9 (52.94%)
BOL-3 (M) 6 (42.86%) 8 (57.14%)
BOL-4 (Spec) 5 (62.50%) 3 (37.50%)
BOL-5 (Spec) 4 (30.77%) 9 (69.23%)
Total 27 (40.91%) 39 (59.09%)
IV-D Overall Performance
The overall performance across all locations is summarized in Table V. A total of 175 output indications were evaluated, with 83 incorrect (47.43%), and 92 correct (52.57%). This indicates that, while improvements are certainly required, the performance of the LLM is promising, also considering the complexity of the different scenarios and paths. While the best performance is obtained for the “simplest” map, i.e., Bologna Airport, challenging scenarios like Orio Center do not appear to affect the performance strongly. Instead, for a case like Bergamo, which also has a simple layout with a single corridor, performances are significantly lower. We argue that this is due to the high number of points of interest and more icons, which appear to overwhelm the LLMs. Our results suggest that it is not the layout or the complexity of the path that affects the performance, but rather the amount of information reported on the map. This can be improved by providing different, simplified views of a map to be used as input to the LLM to facilitate the task. Another phenomenon that we observed, which we can relate to the map’s visual appearance, is the types of errors made for the Orio Center case. In some situations, the LLM gave indications of routes that were not physically possible due to the presence of walls, which were not clearly recognizable due to the similar shades of colours on the map. We thus argue that an image with more contrasting colours could lead to better performance. This can be done in a pre-processing phase, together with removing the legend. It should be highlighted that the maps were all coloured, and in future works, we plan to evaluate the performance considering high-contrast greyscale images, which could simplify the task by discarding irrelevant colour information.
Another phenomenon we observed is that, in many cases, such as the one exemplified in Section III, the LLM seems to struggle with adopting the user’s perspective. Instead of providing directions from the point of view of a subject within the corridors, the output appears to be based on the perspective of an external observer looking at the map. These issues could, in principle, be avoided by identifying the user’s viewpoint and position, which can be achieved by asking them to describe what they see to their left, right, and in front of them.
TABLE V: Overall Results
Location Incorrect Correct
Bergamo 31 (60.78%) 20 (39.22%)
Orio Center 25 (43.10%) 33 (56.90%)
Bologna 27 (40.91%) 39 (59.09%)
Total 83 (47.43%) 92 (52.57%)
IV-E Comprehensive Evaluation of Incorrect Answer
In the comprehensive evaluation of the responses generated by the model during testing, it was discovered that 47.43% were incorrect. To understand the nature of these inaccuracies, a detailed analysis was undertaken. It became clear that these incorrect responses were not in accordance with the expected guidelines. However, this analysis also yielded valuable insights, revealing that the guidance provided does not negatively affect the user’s overall direction and progress despite the errors.
An example, taken from the response related to Figure  3, of insights categorized in this analysis as ’Partially Correct’ includes:
  • “… which will also be on your left.” where the guidance is incorrect, but it is not crucial and does not harm the user’s journey. It is merely a wrong choice of reference point along the path. Errors like this can be classified as less severe than others.
  • “and you will see Clayton” Where there was a mistake in the order of two points of interest along the route. According to the directions, the user should see Dainese first and then Clayton, but in this case, the model reversed their order. This, too, is a minor error that does not compromise the overall direction of the user.
The set of these indications considered partially correct constitutes 16% of the total incorrect responses.
In many cases, as seen in this example, there is a sense that the LLM fails to put itself in the user’s perspective (even when specified). The output does not refer to the point of view of a virtual subject within the corridors but gives directions based on the perspective of an observer looking at the map. These issues can, in principle, be prevented by identifying the position of users, which can be done by asking them to describe what they see on their left/right and in front of them.
V Conclusions
This paper evaluates the performance of ChatGPT in providing human-like indications to users in indoor environments represented through visual maps. We considered three different real-world map layouts with different characteristics, and five types of engineered user prompts of different complexity levels. Our preliminary results are promising, with an average of 47.43% correct indications, and a maximum of 52.57%. Furthermore, we noticed that the complexity of the user query and of the layout of the map do not appear to strongly influence the performance. Instead, these appear to be affected by the visual characteristics of the maps, e.g., the degree of contrast between colours and size and the information overload due to the density of points of interest. In future work, we plan to consolidate and improve these results with more systematic tests, including pre-processed maps, e.g., with increasing contrasts and simplified information. Furthermore, to go towards a more practical application of the envisioned approach, we plan to pre-process natural user queries so that these can be translated into engineered user prompts, similar to the ones used in our tests. These improvements aim to make the system more robust and user-friendly, paving the way for practical applications in real-world indoor navigation systems.